UARS Weekly Status Report

24 November 2000




********************************************************************

                UARS WEEKLY STATUS REPORT - 24 NOV 2000

********************************************************************

***********
SOLSTICE
***********

Nothing received.

***********
SUSIM
***********

Nothing received.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

Nothing received.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

On 21 November the MPG participated in the UARS Status Meeting
which was held to report the status of efforts to use data from
the three-axis magnetometer (TAM) for attitude control of the
UARS spacecraft.  At this meeting it was reported that
development and testing of flight software is nearing completion. 
Tentative plans, assuming no new problems are discovered during
further testing, are to uplink the software changes to the on-
board computer during the week of 4-8 December 2000, and then
begin on-board processing of TAM data in "monitor" mode. 
(Attitude control of UARS will continue to use the Fine Sun
Sensor, backed up by ground processed TAM data, until reliable
operation of the on-board TAM data processing is verified.)

Daily processing of instrument Daily Activity Plans continues 
with no problems.  The MPG continues to coordinate with the 
Flight Operations Team and to assist the Project Scientist 
with the scheduling of instrument operations.  The published
instrument operations schedule is complete to 3 January 2001. 

***********
CDHF
***********

STATUS                          
The CDHF processed current day data through level 3 for all active
   instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
   (C) = Complete, (F)= Failed  (L) = Days remaining to reprocess.
            2 days SEN (5.0) PROD (1992)      (2R-2C=0L)
DAAC DATA:
   -  Some DAAC MLS V5 data has been transferred to the DAAC.  
   -  WINDII requested Version 11 be sent to the DAAC from start of mission to
      31-Aug-1997.  WINDII is in the process of updating the UARS_QUALITY
      field which is a DAAC requirement.            
UARS CDHF REHOST:
   -  Need positive response from all PI End Users prior to Rehost transition.
   -  PI validation of output products and concerns are being addressed.
                  
***********
FOT
***********
                          REPORTING PERIOD

This report covers 11 November 2000 (Orbit 50116, GMT Day 316) through 17 
November 2000 (Orbit 50220, GMT Day 322).

                       SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:

       
                    INSTRUMENT OPERATIONAL CHANGES


Instrument      Time           Orbit           Comment
SUSIM           321/1729       50201           Resume Ops
PEM             322/1406       50214           AXIS ON
WINDII          322/1949       50218           Power ON


The beta angle ranged between the angles of 76.1 degrees to 60.8 degrees for 
this report period.  The beta angle is now decreasing towards a minimum angle 
of 2 degrees on 3 December 2000 (DOY 337).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 1.2 and 0.9 degrees C respectively.  Additional battery performance data 
for each battery on 10 November 2000 and 17 November 2000 are provided in 
Appendix D.

The clock error ranged from 11.0 to 10.8 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

           OUT-OF-LIMITS  =  02
           ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

           FOT           =  15
           SUSIM         =  01
           PEM           =  01
           WINDII        =  01

UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

           FOT           =  08

One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/221  -  2287.496937
 Transponder B:  00/305  - 2287.499087
 
 
 
                       GROUND SYSTEM OPERATION
 
 Eight (8) operational support problems occurred this period.  Anomaly Reports 
are included in Appendix E.
 
The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 00/321 show the observatory data loss to be 301 hours, 06 
minutes, 51.616 seconds (an increase of 01 hours, 46 minutes and 35.952 seconds 
since the last report period).  This is a 0.39 percent data loss which equals a 
99.61 percent data capture for the mission.
 
The increase of 01 hours, 46 minutes and 35.952 seconds occurred over three (3) 
days:
 
 0 minutes, 24.400 seconds on Day 315 due to a data hit
 49 minutes, 25.744 seconds on Day 316 due to a S/C ATC halt TTR# 23135
 51 minutes, 42.896 seconds on Day 317 due to Interface card problems at 
  White Sands
 01 minutes, 18.512 seconds on Day 318 due to Late acquisition TTR#14170
 03 minutes, 30.304 seconds on Day 319 due to Late acquisition TTR#23157
  0 minutes, 14.096 seconds on Day 321 due to a data hit.

                        OTHER SIGNIFICANT EVENTS

                        UARS SATELLITE OPERATIONS

On 07 November 2000 a +10 degree yaw offset maneuver was performed.  This was 
done in order to maintain sun presence over the course of the high beta period.
On November 11, it became necessary to yaw the spacecraft an additional degree 
in the positive direction making the total an 11 degree yaw in order to keep 
the sun in the field of view of the FSS.  The spacecraft was then yawed back to 
its previous +10 degree yaw attitude on November 13 where it remained until the 
November 14th yaw which would bring the spacecraft back to a 0 degree yaw 
offset.

UARS also experienced an Absolute Time Command Processor halt on 11 November 
2000.  This anomaly was due to the fact that the second segment of the ATC was 
not loaded.  Since all of the transmitter on commands are stored in the ATC the 
subsequent pass was missed and the data lost was nonrecoverable.  The last pass 
taken before the halt was a non-coherent pass so the commands to put the 
spacecraft into coherent mode were also sent manually before nominal operations 
could continue.

Testing continues daily to get the TAM and FSS attitude control solution 
operational.  The UARS testing and training simulator (UTTS) is being used for 
all testing.  In order to insure the validity of the test the FOT has been 
debugging some problems with the simulator attitude models.  

                                 MEETINGS

The FOT attended a SERB meeting on November 15, 2000 at 2:00 pm.


                         FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for December 02, 2000 (DOY 337).  This 
maneuver will be a Reverse to Forward Flight maneuver (Northern Hemisphere 
viewing).

Attachments:  Appendix A,  Real-time Commands Sheets
              Appendix B,  Out-Of-Limits & ALERT Occurrences
              Appendix C,  Clock & Solar Array Rate Corrections
              Appendix D,  UARS Battery Performance Data
              Appendix E,  UARS Anomaly Reports

                                APPENDIX A
                         REAL-TIME COMMAND SHEETS
                        Week ending 17 November 2000

ORBIT     AOS         NO.        SUBSYS     REASON
-----     ---         ---        ------     ------
50117     316/0236    00-636     OBC        JUMP ATC LOCATION
50118     316/0342    00-637     OBC        XMTR NOOPS D-316
50118     316/0342    00-638     CDH        SWITCH TO COHERENT MODE
50120     316/0741    00-639     OBC        Clk Adj from 41.2 to 20.6 msec/day
50120     316/0741    00-640     OBC        GENASCII TABLE 51
50121     316/0923    00-641     OBC        GENASCII TABLE 51 (YAW -1 DEG)
50133     317/0419    00-642     CDH        SWITCH TO COHERENT MODE
50142     317/1806    00-643     OBC        Clk Adj from 20.6 to 40.1 msec/day
50155     318/1542    00-644     MACS       YAW FROM 11 DEG TO 10 DEG
50179     320/0524    00-645     NBTR       IN RECORD
50180     320/0702    00-645     NBTR       STOP - HI TEMP
50181     320/0848    00-645     NBTR       PLAYBACK
50186     320/1630    00-1630    OBC        Clk Adj from 41.2 to 20.6 msec/day
50200     321/1548    00-647     OBC/ACS    YAW BACK TO 0 DEG OFFSET
50201     321/1631    00-647     OBC/ACS    LOAD NOMINAL RATE ERROR LIMITS
50201     321/1729    00-648     SUSIM      RESUME OPS
50214     322/1406    00-649     PEM        AXIS ON
50218     322/1949    00-650     WINDII     Power ON
_______________________________________________________________________________

                                  APPENDIX B
                      OUT-OF-LIMITS & ALERTS OCCURRENCES
                         Week ending 17 November 2000

The following Out-Of-Limits occurrences were detected:

MNEMONIC           STATE           DESCRIPTION          ORBITS
--------           -----           -----------          ------
NBTUTEMP2           Y-LO           NBTR TU TEMP         50179-181           
NBTUTEMP2           R-LO           NBTR TU TEMP         50180           


The following ALERT occurrences were detected:

MNEMONIC           DESCRIPTION                          ORBITS
--------           -----------                          ------
                   -None this report period

_______________________________________________________________________________

                                  APPENDIX C
                    CLOCK & SOLAR ARRAY RATE CORRECTIONS
                          Week ending 17 November 2000

                         CLOCK DRIFT RATE ADJUSTMENTS

ORBIT           TIME               CLOCK RATE CHANGE
-----           -------            ---------------------------------
50120           316/0744           Clk Adj from 41.2 to 20.6 msec/day
50142           317/1820           Clk Adj from 20.6 to 41.2 msec/day
50186           320/1634           Clk Adj from 41.2 to 20.6 msec/day

_______________________________________________________________________________

                                  APPENDIX D
                       UARS BATTERY PERFORMANCE DATA
                        Week ending 17 November 2000


           10 Nov 2000 - GMT Day 315 Beta = 76.1 deg, SA Pos = 269 deg
      DIFF V (mV)     TEMP (C)      CURR (amp)   EON V    AVE C/D   AVE
BAT   MAX/MIN         MAX/MIN       MAX/MIN      MIN      RATIO     DOD(%)
---   -----------     -------       ----------   ------   -------   ------
1     +728.0/+728.0   -2.73/-3.05   +0.0/+0.0    14.2     0.000     0.0
2     +11.2/0.0       +2.60/+2.60   +15.6/+9.6   31.7     1.040     13.6
3     0.0/-28.0       +1.97/+1.97   +09.6/+9.2   31.7     1.040     13.1


           17 Nov 2000 - GMT Day 322 Beta = 60.8 deg, SA Pos = 269 deg

       DIFF V (mV)     TEMP (C)     CURR (amp)    EON V   AVE C/D   AVE
BAT    MAX/MIN         MAX/MIN      MAX/MIN       MIN     RATIO     DOD(%)
---    ------------    --------     ----------    ------  -------   ------
1      +728.0/+728.0   -3.05/-4.31  +0.0/+0.0     14.2    0.000     0.0
2      +11.2/-22.4     +3.87/+2.60  +23.2/-14.4   26.7    1.030     16.8
3      +11.2/-33.6     +2.60/+1.66  +24.4/-13.6   26.7    1.020     16.4


           V/T LEVEL CHANGES
ORB#           TIME                   LVL FR-TO
----           ----                   ---------
               -no changes             V/T 5
_______________________________________________________________________________

 APPENDIX E
                           UARS ANOMALY REPORTS
                        Week ending 17 November 2000

                      NEW ANOMALY REPORTS GENERATED
AR NO.    SUBSYS      ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------    ------      -----     ----        ---     -----------------
00-050    FOT         50116     316/0050Z   23135   ATC HALT & S/C MISCONFIG
00-051    FOT         50131     317/0053    23142   TDRSS-E down: 10:42 D/L
00-051    FOT         50132     317/0238    23142   TDRSS-E down: 10:42 D/L 
00-052    FOT         50133     317/0419    23142   WSGT ground system prob.
00-047    FOT         50147     318/0240    14170   Generic Late Acquisition
00-053    FOT         50168     319/1113    23157   Late Acquisition
00-047    FOT         50185     320/1546    14170   Generic Late Acquisition  
00-047    FOT         50206     322/0027    14170   Generic Late Acquisition

                   ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.    SUBSYS      ORBIT     TIME         TTR     PROBLEM/DATA LOSS
------    ------      -----     ----         ---     -----------------
00-048    FOT         50025     309/2255     N/A     ATC HALT
00-049    FOT         50051     311/1623     N/A     HGA at S/W stop limit
00-050    FOT         50116     316/0050Z    23135   ATC HALT & S/C MISCONFIG
00-051    FOT         50131     317/0053     23142   TDRSS-E down: 10:42 D/L
00-052    FOT         50133     317/0419     23142   WSGT ground system prob.


                   ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.    SUBSYS      ORBIT     TIME         TTR      PROBLEM/DATA LOSS
------    ------      -----     ----         ---     -----------------
           -None this report period

                           OPEN ANOMALY REPORTS
AR NO.    SUBSYS      ORBIT     TIME        TTR       PROBLEM/DATA LOSS
------    ------      -----     ----        ---       -----------------
00-047    FOT         see list  Nov 00      14170     Generic Late Acquisition-
053       FOT         50168     319/1113    23157     Late Acquisition


	 	SEVNTFW00/46